# Performance in solving the Hermitian and pseudo-Hermitian Bethe-Salpeter equation with the Yambo code

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics Codebases.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
The Bethe-Salpeter Equation (BSE) is the state-of-the-art first-principles approach to compute neutral excitations in material science [1, 2]. It can be formulated as a non-Hermitian structured eigenvalue problem given by
@@
We refer to the full BSE matrix as $H^{2p}$, and to its structure as pseudo-Hermitian [3, 4, 5, 6]. A matrix is said to be pseudo-Hermitian, if there exists a Hermitian linear operator $\eta$, such that $H^{\dagger}=\eta H\eta^{-1}$ [7]. In the case of BSE, the matrix $H^{2p}$ satisfies the condition with respect to the $\eta=\Omega$, and is thus $\Omega$-pseudo-Hermitian. The BSE Hamiltonian, as shown in Eq. (3), can be expressed as the matrix product of two Hermitian matrices. If the $M$ matrix is positive definite, which can be verified by Cholesky factorization, then the BSE problem reduces to a Hermitian-definite generalized eigenvalue problem with real eigenvalues [3]. Otherwise, it should be treated as a general non-Hermitian eigenvalue problem. Additionally, it can be shown that BSE matrix has a block containing transitions of positive energy (resonant matrix $R$), a block containing transitions of negative energy (anti-resonant matrix $A$), and two blocks which mix positive and negative energy transitions (coupling matrices $C$ and $C^{\dagger}$). Neglecting the coupling term, i.e., setting $C=0$, corresponds to the Tamm-Dancoff approximation (TDA) [1, 8]. Regardless of $M$ being positive definite or not, the BSE matrix in Eq. (1) must have a real eigenspectrum to describe a physical system, provided that lifetimes effects are not included, i.e., all the matrices are frequency independent, as it is usually the case. The structure of the BSE also allows further simplifications of the eigenvalue problem in the special case where R = A, as discussed later. This enables more efficient solution methods that go beyond the standard Hermitian-definite generalized eigenvalue approach [9].
@@
The BSE eigenvalue problem is coded in many condensed-matter (and quantum chemistry) packages, both open-source, such as Yambo [10, 11], BerkeleyGW [2, 12], Abinit [13, 14], Exc [15], Exciting [16, 6, 17], GPAW [18, 19], Simple [20], MolGW [21] and commercial or private, such as VASP [22, 5], TurboMole [23], and Fiesta  [24, 25]. Considering also TD-DFT and RPA, the number of scientific codes implementing the eigenvalue problem of Eq. (1) becomes very large, including applications in quantum chemistry and nuclear physics. Most of them use the Hermitian-definite generalized eigenvalue solvers implemented in eigensolver packages. In condensed matter, most BSE applications are done within the TDA, where the BSE eigenvalue problem becomes Hermitian. Indeed, at variance with the case of isolated systems studied in quantum chemistry, the BSE can be recast into a Hermitian quadratic problem only in presence of time-reversal symmetry [5]. However, there are physical cases where going beyond the TDA is needed, and time-reversal is not available [26]. Thus, the full pseudo-Hermitian problem must be addressed [3, 4, 6].
For the case of optical properties, we have $A=R^{*}$, while $C=C^{T}$. The size of both $R$ and $C$ is $N\times N$. Defining $N_{\mathbf{k}}$ as the number of k-points in the full Brillouin zone, $N_{v}$ as the number of included occupied (valence) states, and $N_{c}$ as the number of unoccupied (conduction) states, $N=N_{\mathbf{k}}\,N_{v}\,N_{c}$ is the total number of optical transitions. Numerically, the most demanding step is the building of the $R$ and $C$ matrices. This operation scales as $\mathcal{O}(N^{2})$, with a large prefactor that depends on the square of the basis-set size. For plane-wave codes, this corresponds to $\mathcal{O}(N_{G}^{2})$ where $N_{G}$ is the number of G-vectors. This is why, with the advent of HPC machines and complex structures, a significant effort has been devoted to improve the matrix generation algorithm, taking advantage of MPI, OpenMP and GPU porting schemes such as CUDA Fortran, OpenMP5 or OpenACC. Thanks to these developments, matrices with a size of the order $N\propto 10^{5}$ can be easily generated. Even larger matrices can be generated by means of interpolation techniques [27, 28].
@@
In the present work, we analyze the performances of BSE solvers, two based on direct diagonalization, and the other based on an iterative scheme, and we use them to solve prototype BSE matrices. The two approaches have been implemented in the Yambo code, a plane-wave code, taking advantage of the interfaces with libraries such as ScaLAPACK [29], ELPA [30, 31, 32], PETSc [33] and SLEPc [34]. We consider both the Hermitian BSE within the TDA, and the full structured pseudo-Hermitian BSE matrix. For the latter case, we discuss in detail how we take advantage of the pseudo-Hermitian structure in both solvers to reduce the computational time. We provide parallel performances, using a local cluster for CPU only simulations, and the Leonardo machine for GPU ported simulations. An analysis of the memory used on the host is also provided.
@@ 2.1.1 The Diagonalization Scheme for Full BSE Matrix (Coupling Case)
We briefly outline the special algorithm for diagonalizing the BSE Hamiltonian beyond TDA given in Eq. (1) when $A=R^{*}$ and $C=C^{T}$, as shown in Ref. [9]. The first step in solving the BSE Hamiltonian, $H^{2p}$, with this special structure is to construct a real symmetric matrix $M_{r}$ of the same dimension, which satisfies the relation $Q^{\dagger}H^{2p}Q=-iJM_{r}$, where $J$ is a real skew-symmetric matrix and $Q$ is a unitary matrix given by:
@@
If $M$ is positive definite, then $M_{r}$ is also positive definite [9], and we can perform a Cholesky decomposition on the real symmetric matrix $M_{r}$, i.e., $M_{r}=LL^{T}$ where $L$ is a real lower triangular matrix. This allows us to cast the original eigenvalue problem into the following form:
$-iWL^{T}\begin{pmatrix}\bar{X}_{I}\\ \bar{Y}_{I}\end{pmatrix}=\omega_{I}L^{T}\begin{pmatrix}\bar{X}_{I}\\ \bar{Y}_{I}\end{pmatrix}$ (6)
where $W=L^{T}JL$ is a real skew-symmetric matrix, and $-iW$ is Hermitian, having the same eigenvalues as the original BSE Hamiltonian. In Eq. (6), the eigenvalues of $W$ differ from the $H^{2p}$ eigenvalues by a factor of $-i$. The eigenvalues of a skew-symmetric matrix come in pairs and are purely imaginary, the eigenvalues of $H^{2p}$ are real and occur in pairs, i.e., $(-\omega_{I},\omega_{I})$. Therefore, the entire BSE problem can be viewed as a real skew-symmetric eigenvalue problem.
In diagonalization, the tridiagonalization step is the computationally expensive part. The algorithm described here enables the tridiagonalization of a real matrix instead of a Hermitian one, significantly reducing both floating-point operations and storage requirements. Once the tridiagonal form of the skew-symmetric matrix is obtained, multiplying the result by $-i$ allows the use of Hermitian tridiagonal solvers to compute the eigenvectors of $W$ [32, 35]. After computing the eigenvectors, a back-transformation is applied to retrieve those of the BSE matrix. Moreover, the left eigenvectors for positive eigenvalues, and the left/right eigenvectors corresponding to its negative partner can be determined without explicit computation [9]. More details on the implementation can be found in Ref. [36]. Consequently, we only need to obtain the right eigenvectors associated with positive eigenvalues. The left and right eigenvectors of the positive and negative eigenvalues for the BSE matrix are related as follows [9]:
$\begin{pmatrix}\text{Right eigenvector for }\omega_{I}&\text{Left eigenvector for }\omega_{I}\\ \begin{pmatrix}X_{I}\\ Y_{I}\end{pmatrix}&\begin{pmatrix}X_{I}\\ -Y_{I}\end{pmatrix}\\ \text{Right eigenvector for }-\omega_{I}&\text{Left eigenvector for }-\omega_{I}\\ \begin{pmatrix}Y_{I}^{*}\\ X_{I}^{*}\end{pmatrix}&\begin{pmatrix}-Y_{I}^{*}\\ X_{I}^{*}\end{pmatrix}\end{pmatrix}$ (7)
@@
Unlike Hermitian-definite generalized eigenvalue solvers, this method reduces floating-point operations for the most computationally demanding steps. Furthermore, when a direct Hermitian-definite generalized eigenvalue solver is used on the BSE matrix, it can destroy the special properties of this matrix and potentially break degeneracies [9], which is highly undesirable in these types of calculations. Aside from improving efficiency, this approach is also robust to such errors.
@@ 2.1.2 Interface with ScaLAPACK and ELPA
To efficiently diagonalize the excitonic Hamiltonian in Yambo, both in the Hermitian case and, for the coupling case, using the algorithm described above, an MPI redistribution from the Yambo parallel structure to a block-cyclic layout is performed in two main steps. First, the matrix elements are locally grouped along with their corresponding global indices based on their destination rank. Then, a single MPI_AlltoAllv operation is used to transfer the data to their respective destination according to the block-cyclic scheme. Once the matrix is distributed correctly, standard functions from linear algebra libraries can be applied for efficient computations. Here we use ScaLAPACK [37] and ELPA. For the TDA case, i.e., when the BSE matrix is Hermitian, we use standard Hermitian eigensolvers. ELPA provides two types of Hermitian solver, with type-2 set as the default in Yambo. In ScaLAPACK, the p?heevr solver is used. Both solvers allow one to compute the lowest part of the spectrum. One advantage of ELPA is that its newer versions are specifically optimized for skew-symmetric matrices [32]. This makes ELPA particularly well-suited for solving pseudo-Hermitian or skew-symmetric eigenvalue problems. Implementing such solvers with ELPA is more straightforward, whereas using ScaLAPACK for the same purpose would require additional modifications to handle the skew-symmetric structure. However, ScaLAPACK enables the extraction of eigenvectors within a specified range of eigenvalues or indices.
The interface described above has recently been implemented through the development of a new library [36], which is interfaced via the Yambo code. With this new implementation of the diagonalization solver, capable of computing either the full or partial eigenspectrum, we address the long-standing bottleneck related to the diagonalization of BSE matrices in the Yambo code. Although the code is not yet available in the official Yambo release, it is freely distributed through a recently created fork of the code called Lumen [38, 39].
@@ 2.2.1 The SLEPc Solver for the Full BSE Matrix (Coupling Case)
Prior to the release of SLEPc 3.22, the eigenproblem was solved with a standard non-Hermitian two-sided eigensolver, with the possibility to use shift-and-invert to accelerate the convergence of the desired eigenvalues, which are those closest to zero, and lie in the innermost part of the spectrum because of its symmetry. Since the release of SLEPc 3.22, three new methods are available based on [40, 41], which are three variations of a structure-preserving Lanczos iteration with a thick restart.
@@
The Shao method does not require the use of the computationally expensive shift-and-invert technique because it computes the eigenvalues of $H^{2p}$, folding the spectrum so that the internal eigenvalues become external and converge first. A last step is required to obtain the positive eigenvalues and associated eigenvectors of $H^{2p}$. The negative eigenvalues and their eigenvectors, and the left eigenvectors, can be obtained trivially from them. The details of the SLEPc solvers for the Bethe–Salpeter eigenproblem are available in [42].
@@ 2.2.2 MPI Distribution and GPU Porting in SLEPc
In the interface between Yambo and SLEPc two different schemes are available. The first, or shell approach, is based on shell matrices. These are matrices that do not store their values explicitly but instead define how operations on them are carried out. This allows the solver to reuse the existing Yambo data and parallel structures directly. Because Yambo efficiently stores the BSE matrix in memory by exploiting its symmetry properties, this approach requires less memory. However, it comes at the cost of relying on a parallel scheme that is optimized for building of the BSE matrix rather than for solving the eigenvalue problem. Additionally, this approach does not support GPU acceleration. The second approach, called the explicit approach is based on a redistribution of the BSE matrix from the block parallel Yambo structure to the SLEPc structure. In SLEPc, MPI parallelization is based on standard one-dimensional row-wise partitioning of matrices and vectors. This is the scheme used by PETSc, which is the library on which SLEPc is built. In the case of GPU simulations, the underlying local computations are offloaded to the NVIDIA GPU using CUDA (or HIP for AMD GPUs). The GPU porting of the Yambo-SLEPc interface was achieved by defining the redistributed matrix directly on GPU, specifically using the SLEPc matrix type MATDENSECUDA. As in the diagonalization case, this new algorithm is not yet included in the official Yambo release, but it is freely available through a recently developed fork of the code called Lumen [38].

### v3

# Solvers for the Hermitian and the pseudo-Hermitian Bethe-Salpeter equation in the Yambo code: Implementation and Performance

@@ Abstract
keywords:
Bethe–Salpeter equation, matrix diagonalization, optical properties, pseudo-Hermitian eigenvalue problem, iterative schemes, Krylov-Schur
journal: Computational Materials Science
[…]
@@ 1 Introduction
The Bethe-Salpeter Equation (BSE) is the state-of-the-art first-principles approach to compute neutral excitations in material science [Onida2002, Rohlfing2000]. It explicitly accounts for electron–hole interactions, allowing for an accurate description of excitonic effects and thereby enabling reliable prediction of optical properties in a wide range of materials, including 2D materials [2d_materials_1, 2d_materials_2], wide band-gap insulators and transition metal oxides [wide_gap_insulators, transition_metal_oxides], semi-conductors, and magnetic materials [semiconductors, magnetic_materials]. It was also recently used to compute magnons in magnetic materials [magnon_in_cri3]. It can be formulated as a non-Hermitian structured eigenvalue problem given by
@@
We refer to the full BSE matrix as $H^{2p}$, and to its structure as pseudo-Hermitian [Gruuning2009, Gruening2011, Sandre2015, Vorwerk2019]. A matrix is said to be pseudo-Hermitian, if there exists a Hermitian linear operator $\eta$, such that $H^{\dagger}=\eta H\eta^{-1}$ [Mostafazadeh_2002]. In the case of BSE, the matrix $H^{2p}$ satisfies the condition with respect to the $\eta=\Omega$, and is thus $\Omega$-pseudo-Hermitian. The BSE Hamiltonian, as shown in Eq. (3), can be expressed as the matrix product of two Hermitian matrices. If the $M$ matrix is positive definite, which can be verified by Cholesky factorization, then the BSE problem reduces to a Hermitian-definite generalized eigenvalue problem with real eigenvalues [Gruuning2009]. Otherwise, it should be treated as a general non-Hermitian eigenvalue problem. Additionally, it can be shown that BSE matrix has a block containing transitions of positive energy (resonant matrix $R$), a block containing transitions of negative energy (anti-resonant matrix $A$), and two blocks which mix positive and negative energy transitions (coupling matrices $C$ and $C^{\dagger}$). Neglecting the coupling term, i.e., setting $C=0$, corresponds to the Tamm-Dancoff approximation (TDA) [Onida2002, Kammerlander_2012]. Regardless of $M$ being positive definite or not, the BSE matrix in Eq. (1) must have a real eigenspectrum to describe a physical system, provided that lifetimes effects are not included, i.e., all the matrices are frequency independent, as it is usually the case. The structure of the BSE also allows further simplifications of the eigenvalue problem in the special case where $R=A$, as discussed later. This enables more efficient solution methods that go beyond the standard Hermitian-definite generalized eigenvalue approach [SHAO2016148].
@@
The BSE eigenvalue problem is coded in many condensed-matter (and quantum chemistry) packages, both open-source, such as Yambo [Marini2009, Sangalli2019], BerkeleyGW [Rohlfing2000, Deslippe2012], Abinit [Gonze2016, Gonze2020], Exc [Gatti2013], Exciting [Gulans2014, Vorwerk2019, Hennekek2020], GPAW [Mortense2024, Thygesen2017], Simple [Prandini2019], MolGW [Bruneval2016] and commercial or private, such as VASP [Kresse1999, Sandre2015], TurboMole [Krause2017], and Fiesta  [Faber2014, Knysh2023]. Considering also TD-DFT and RPA, the number of scientific codes implementing the eigenvalue problem of Eq. (1) becomes very large, including applications in quantum chemistry and nuclear physics. Most of them use the Hermitian-definite generalized eigenvalue solvers implemented in eigensolver packages. In condensed matter, most BSE applications are done within the TDA, where the BSE eigenvalue problem becomes Hermitian. Indeed, at variance with the case of isolated systems studied in quantum chemistry, the BSE can be recast into a Hermitian eigenproblem only in presence of time-reversal symmetry [Sandre2015]. However, there are physical cases where going beyond the TDA is needed, and time-reversal is not available [footnote1]. Thus, the full pseudo-Hermitian problem must be addressed [Gruuning2009, Gruening2011, Vorwerk2019].
For the case of optical properties, we have $A=R^{*}$, while $C=C^{T}$. The size of both $R$ and $C$ is $N\times N$. Defining $N_{\mathbf{k}}$ as the number of k-points in the full Brillouin zone, $N_{v}$ as the number of included occupied (valence) states, and $N_{c}$ as the number of unoccupied (conduction) states, $N=N_{\mathbf{k}}\,N_{v}\,N_{c}$ is the total number of optical transitions. Numerically, the most demanding step is the building of the $R$ and $C$ matrices. This operation scales as $\mathcal{O}(N^{2})$, with a large prefactor that depends on the square of the basis-set size. For plane-wave codes, this corresponds to $\mathcal{O}(N_{G}^{2})$ where $N_{G}$ is the number of G-vectors. This is why, with the advent of HPC machines and complex structures, a significant effort has been devoted to improve the matrix generation algorithm, taking advantage of MPI, OpenMP and GPU porting schemes such as CUDA Fortran, OpenMP5 or OpenACC. Thanks to these developments, matrices with a size of the order $N\propto 10^{5}$ can be easily generated. Even larger matrices can be generated by means of interpolation techniques [Gillet2016, Alliati2022].
@@
The size of the Bethe–Salpeter Equation (BSE) matrix plays an important role in the study of excitons and the simulation of optical properties. In general, using the full BSE matrix with as many bands as possible is desirable in order to properly converge excitonic energies and to capture exciton physics within a material. More specifically, matrices with size of the order $10^{4}$ - $10^{5}$ can be easily reached when studying 2D heterostructures [PhysRevB.95.075423, PhysRevB.110.035118], systems with defects [PhysRevB.109.085127, dg13-y4kj], systems that involve molecular interactions such as functionalized materials (e.g., functionalized graphene), or systems containing impurities. In some cases, neglecting the coupling term is not an option. For instance, the coupling term is known to be essential when studying localized states which are important in some of the applications mentioned above. Moreover, when moving to large-scale systems, such as supercells or nanostructures, the BSE matrix becomes extremely large. In such cases, efficient diagonalization algorithms are necessary to make the calculations computationally feasible.
In the present work, we analyze the performances of BSE solvers, two based on direct diagonalization, and the other based on an iterative scheme, and we use them to solve prototype BSE matrices. The two approaches have been implemented in the Yambo code, a plane-wave code, taking advantage of the interfaces with libraries such as ScaLAPACK [Blackford1997], ELPA [Auckenthaler2011, Marek2014, penke2020], PETSc [petsc-ug] and SLEPc [slepc-toms]. We consider both the Hermitian BSE within the TDA, and the full structured pseudo-Hermitian BSE matrix. For the latter case, we discuss in detail how we take advantage of the pseudo-Hermitian structure in both solvers to reduce the computational time. We provide parallel performances, using a local cluster for CPU only simulations, and the Leonardo machine for GPU ported simulations. An analysis of the memory used on the host is also provided.
@@ 2.1.1 The Diagonalization Scheme for Full BSE Matrix (Coupling Case)
We briefly outline the special algorithm for diagonalizing the BSE Hamiltonian beyond TDA given in Eq. (1) when $A=R^{*}$ and $C=C^{T}$, as shown in Ref. [SHAO2016148]. The first step in solving the BSE Hamiltonian, $H^{2p}$, with this special structure is to construct a real symmetric matrix $M_{r}$ of the same dimension, which satisfies the relation $Q^{\dagger}H^{2p}Q=-iJM_{r}$, where $J$ is a real skew-symmetric matrix and $Q$ is a unitary matrix given by:
@@
If $M$ is positive definite, then $M_{r}$ is also positive definite [SHAO2016148], and we can perform a Cholesky decomposition on the real symmetric matrix $M_{r}$, i.e., $M_{r}=LL^{T}$ where $L$ is a real lower triangular matrix. This allows us to cast the original eigenvalue problem into the following form:
$iWL^{T}\begin{pmatrix}\bar{X}_{I}\\ \bar{Y}_{I}\end{pmatrix}=\omega_{I}L^{T}\begin{pmatrix}\bar{X}_{I}\\ \bar{Y}_{I}\end{pmatrix}$ (6)
where $W=L^{T}JL$ is a real skew-symmetric matrix, and iW-iW is Hermitian, having the same eigenvalues as the original BSE Hamiltonian. In Eq. (6), the eigenvalues of $W$ differ from the $H^{2p}$ eigenvalues by a factor of i-i. The eigenvalues of a skew-symmetric matrix come in pairs and are purely imaginary, the eigenvalues of $H^{2p}$ are real and occur in pairs, i.e., $(-\omega_{I},\omega_{I})$. Therefore, the entire BSE problem can be viewed as a real skew-symmetric eigenvalue problem.
In diagonalization, the tridiagonalization step is the computationally expensive part. The algorithm described here enables the tridiagonalization of a real matrix instead of a Hermitian one, significantly reducing both floating-point operations and storage requirements. Once the tridiagonal form of the skew-symmetric matrix is obtained, multiplying the result by i-i allows the use of Hermitian tridiagonal solvers to compute the eigenvectors of $W$ [penke2020, footnote2]. After computing the eigenvectors, a back-transformation is applied to retrieve those of the BSE matrix. Moreover, the left eigenvectors for positive eigenvalues, and the left/right eigenvectors corresponding to its negative partner can be determined without explicit computation [SHAO2016148]. More details on the implementation can be found in Ref. [Murali2025]. Consequently, we only need to obtain the right eigenvectors associated with positive eigenvalues. The left and right eigenvectors of the positive and negative eigenvalues for the BSE matrix are related as follows [SHAO2016148]:
[…]
@@
Unlike Hermitian-definite generalized eigenvalue solvers, this method reduces floating-point operations for the most computationally demanding steps. Furthermore, when a direct Hermitian-definite generalized eigenvalue solver is used on the BSE matrix, it can destroy the special properties of this matrix and potentially break degeneracies [SHAO2016148], which is highly undesirable in these types of calculations. Aside from improving efficiency, this approach is also robust to such errors.
@@ 2.1.2 Interface with ScaLAPACK and ELPA
To efficiently diagonalize the excitonic Hamiltonian in Yambo, both in the Hermitian case and, for the coupling case, using the algorithm described above, an MPI redistribution from the Yambo parallel structure to a block-cyclic layout is performed in two main steps. First, the matrix elements are locally grouped along with their corresponding global indices based on their destination rank. Then, a single MPI_AlltoAllv operation is used to transfer the data to their respective destination according to the block-cyclic scheme. Once the matrix is distributed correctly, standard functions from linear algebra libraries can be applied for efficient computations. Here we use ScaLAPACK and ELPA. For the TDA case, i.e., when the BSE matrix is Hermitian, we use standard Hermitian eigensolvers. ELPA provides two types of Hermitian solver, with type-2 set as the default in Yambo. In ScaLAPACK, the p?heevr solver is used. Both solvers allow one to compute the lowest part of the spectrum. One advantage of ELPA is that its newer versions are specifically optimized for skew-symmetric matrices [penke2020]. This makes ELPA particularly well-suited for solving pseudo-Hermitian or skew-symmetric eigenvalue problems. Implementing such solvers with ELPA is more straightforward, whereas using ScaLAPACK for the same purpose would require additional modifications to handle the skew-symmetric structure. However, ScaLAPACK enables the extraction of eigenvectors within a specified range of eigenvalues or indices.
Since ScaLAPACK does not provide a dedicated solver for real skew-symmetric matrices, the tridiagonalization routines designed for symmetric matrices can be slightly modified to accommodate them, as both share similar computational structures. For the present implementation, we employ ScaLAPACK’s Hermitian solver by passing the matrix iW-iW, which converts the real skew-symmetric matrix $W$ into a Hermitian form. In future releases, we plan to modify ScaLAPACK’s real tridiagonalization routines to directly handle the skew-symmetric matrix $W$, thereby eliminating the need for this transformation.
The interface described above has recently been implemented through the development of a new library [Murali2025, Ldiago], which is interfaced via the Yambo code. With this new implementation of the diagonalization solver, capable of computing either the full or partial eigenspectrum, we address the long-standing bottleneck related to the diagonalization of BSE matrices in the Yambo code. Although the code is not yet available in the official Yambo release, it is freely distributed through a recently created fork of the code called Lumen [Lumen, Ldiago].
@@ 2.2.1 The SLEPc Solver for the Full BSE Matrix (Coupling Case)
Prior to the release of SLEPc 3.22, the eigenproblem was solved with a standard non-Hermitian two-sided eigensolver, with the possibility to use shift-and-invert to accelerate the convergence of the desired eigenvalues, which are those closest to zero, and lie in the innermost part of the spectrum because of its symmetry. Since the release of SLEPc 3.22, three new methods are available based on [Shao:2018:SPL, Gruning:2011:ITL], which are three variations of a structure-preserving Lanczos iteration with a thick restart.
@@
The Shao method does not require the use of the computationally expensive shift-and-invert technique because it computes the eigenvalues of $H^{2p}$, folding the spectrum so that the internal eigenvalues become external and converge first. A last step is required to obtain the positive eigenvalues and associated eigenvectors of $H^{2p}$. The negative eigenvalues and their eigenvectors, and the left eigenvectors, can be obtained trivially from them. The details of the SLEPc solvers for the Bethe–Salpeter eigenproblem are available in [Alvarruiz:2025:ITL].
@@ 2.2.2 MPI Distribution and GPU Porting in SLEPc
In the interface between Yambo and SLEPc two different schemes are available. The first, or shell approach, is based on shell matrices. These are matrices that do not store their values explicitly but instead define how operations on them are carried out. This allows the solver to reuse the existing Yambo data and parallel structures directly. Because Yambo efficiently stores the BSE matrix in memory by exploiting its symmetry properties, this approach requires less memory. However, it comes at the cost of relying on a parallel scheme that is optimized for building of the BSE matrix rather than for solving the eigenvalue problem. Additionally, this approach does not support GPU acceleration. The second approach, called the explicit approach is based on a redistribution of the BSE matrix from the block parallel Yambo structure to the SLEPc structure. In SLEPc, MPI parallelization is based on standard one-dimensional row-wise partitioning of matrices and vectors. This is the scheme used by PETSc, which is the library on which SLEPc is built. In the case of GPU simulations, the underlying local computations are offloaded to the NVIDIA GPU using CUDA (or HIP for AMD GPUs). The GPU porting of the Yambo-SLEPc interface was achieved by defining the redistributed matrix directly on GPU, specifically using the SLEPc matrix type MATDENSECUDA. As in the diagonalization case, this new algorithm is not yet included in the official Yambo release, but it is freely available through a recently developed fork of the code called Lumen [Lumen].